OpenStudy (anonymous):

−3z − 3(z − 3) = 3 z= (last one lol)

OpenStudy (anonymous):

1

OpenStudy (lilg132):

-3z - 3z + 9 = 3 -6z = 3 - 9 -6z = -6 z = 1

OpenStudy (anonymous):

-3z-3z+9=3 -6z=-6 z=1

OpenStudy (lilg132):

dont forget to click good answer !!!!!

OpenStudy (stormfire1):

Do you need help understanding the steps? You could do these faster yourself :)

OpenStudy (stormfire1):

rather than typing them out

OpenStudy (lilg132):

lool true

OpenStudy (anonymous):

my brother is a genius... hes not around right now

OpenStudy (anonymous):

z=1